# Views Are My Own, But Also Yours: Benchmarking Theory of Mind using Common Ground

### v1

@@ Abstract
Evaluating the theory of mind (ToM) capabilities of language models (LMs) has recently received much attention. However, many existing benchmarks rely on synthetic data which risks misaligning the resulting experiments with human behavior. We introduce the first ToM dataset based on naturally occurring spoken dialogs, Common-ToM, and show that LMs struggle to demonstrate ToM. We then show that integrating a simple, explicit representation of beliefs improves LM performance on Common-ToM.
@@ 1 Introduction
In this paper, we introduce Common-ToM– a question answering benchmark based on naturally-occurring, spoken dialogs in English. [^1] All data, code, and models will be made public. Common-ToM uses the notion of common ground (CG) for evaluating ToM. The CG (Wilkes-Gibbs and Clark, 1992; Stalnaker, 2002) is a set of beliefs mutually shared by all participants in a conversation. Other benchmarks have developed ToM questions based on the Sally-Anne test (Wimmer and Perner, 1983; Baron-Cohen et al., 1985) which checks an understanding of information accessibility to determine beliefs. We observe that when CG is mismatched between the discourse participants, such as at the time of a question or during the repair of a miscommunication, similarly complex problems for ToM arise.
Our main contributions are: (1) arguing that using synthesized data in arguing about the ToM ability of LMs is not conclusive; (2) releasing a corpus for benchmarking ToM based on naturally occurring spoken conversations; (3) showing that LLMs struggle with our benchmark and a simple explicit architecture performs better.
The paper is organized as follows. First, we review some of the relevant literature (§ 2). We then describe the framework and methods used in creating Common-ToM (§ 3). This is followed by experiments and results in § 4: human performance, using zero-shot and fine-tuned LMs, and using our approach for building a system that explicitly represents beliefs (§ 4.4). Finally, we discuss our key takeaways and conclusions (§ 5).
@@ 2 Related Work
While quite a few corpora annotate author belief (Marie-Catherine de Marneffe, 2019; Saurí and Pustejovsky, 2009), there are relatively few corpora annotated explicitly for CG. Horton and Gerrig (2016) focus on limited aspects of CG. In this paper, we use the comprehensive CG corpus presented by Markowska et al. (2023), which we discuss in section 3.
@@
Though useful, these benchmarks are prone to surface-level cues and spurious correlations which have been exploited by LMs to display illusory ToM (Kosinski, 2023). Some work has been done to produce tools based on human-generated text. Bara et al. (2021), like us, exploit the relationship between CG and ToM: they design an experimental setup in order to record dialogs; the dialogs are motivated by the experimental setting; and they are written dialogs. More recently, Ma et al. (2023) combine the aforementioned datasets (and more) to create a composite benchmark. We build on this work conceptually by producing a new dataset which, in contrast with the work of Bara et al. (2021), is based on dialogs which are collected independently of our interest in ToM; the dialogs are spontaneous and not guided by any experimental setting; and they are spoken dialogs. We note that co-presence (including telepresence) has been identified as an important (though not required) part of human ToM (Galati and Brennan, 2021), and our corpus allows the study of ToM under this common condition using naturalistic data.
@@ 3 Common-ToM
Framework In philosophy, CG is sometimes treated as the mutual beliefs between two agents Stalnaker (2002), separate from their cognitive states, while in cognitive science it is common to model CG as the belief of an agent about what they and another agent mutually believe (Brown-Schmidt and Duff, 2016). We adopt the latter definition as it is allows CG to represent scenarios like false belief, which would be impossible under the former. This also makes the relationship between CG and ToM explicit: If I believe a proposition is in the CG with you, it means by definition that I believe that you also believe it is in CG – thus, it entails a ToM assumption.
We also follow cognitive literature in assuming that humans do not have a full representation of their interlocutor’s cognitive state at all times, but instead can make necessary inferences when needed (Horton and Brennan, 2016). Specifically, CG does not mean that all consequences of CG are present at all times, and performing the inferences from CG (high-degree knowledge questions, for example) takes time and may be errorful. Our experiments (§ 4) with human annotators confirm this assumption, but also show that humans are much better than zero-shot LMs because they are in fact modeling CG, and can answer the higher-order belief questions as needed from their models of CG. Farkas and Bruce (2010); Eckardt (2016) propose deterministic heuristic algorithms for inferring CG from text. All this prior work motivates our neuro-symbolic approach (4.4).
Base Corpus Markowska et al. (2023) use a corpus of 1,710 turns of dyadic dialog across four conversations from CALLHOME, and annotate them for CG. They extract events (i.e., propositions) evoked by each turn and label the beliefs of each discourse participant (DP) towards those events as certainly true (CT+), possibly true (PS), certainly not true (CT-), or they label the DP as not having a belief (NB). Then they determine for each DP if the speaker has just added the event to their view of the CG (JA), already has the event in the CG (IN), or rejects it from the CG (RT). We use NA to indicate this. Note that the annotation assumes that the CG is not actually shared, but rather independently hypothesized by each DP.  [^2] We refer the reader to Appendix A and Markowska et al. (2023) for additional details regarding the corpus.
@@
To motivate the query creation process, consider the examples in Table 1. The proposition under question is “B is now not smoking”. The annotations from the CG corpus for this proposition throughout discourse time (which we simply measure in turns) are shown above and resulting first order (A/B believes) and second order (A/B believes that B/A believes) questions are shown below.
@@
At the time indicated, is it the case that (((A/B believes that)${}^{1}$ B/A believes that)${}^{2}$ A/B believes that)${}^{3}$ it is {certainty} true that {proposition}?
@@
Finally, (3) we determine the answers to the generated queries using a set of rules. For first order questions this is straightforward and we simply check that the belief annotation matches the question with one caveat. If A believes proposition $p$ is certainly true and the query asks if A believes $p$ is possibly true, we consider the answer to be yes. For higher order queries we must use the CG annotations. To illustrate, consider the 2nd order question whether A believes that B believes some proposition $p$. (1) If the certainty is positive polarity (i.e., certainly or possibly) and A believes the position to be CG (i.e., JA or IN), then resolve the answer just as in the first order case described above. (2) Otherwise, if the certainty is negative polarity (i.e. certainly not) and A is aware it was rejected from CG (i.e., RT) and the certainty of the query matches the certainty of B, then label the question yes. (3) For all other cases, label the answer no. The whole set of heuristics are in Appendix E.
@@ 4.1 Setup
Data We train on the CG corpus (Markowska et al., 2023). The exact train/test split is shown in Table 3 in Appendix A.
Random Performance We perform a random baseline by randomly selecting “Yes” or “No” answers with their probabilities proportional to their respective frequencies in the training set.
Human Performance We randomly sample sixty queries from our test set with twenty questions from each order. Annotators are graduate students given instructions on the task.
@@ 4.2 Zero-Shot
We use the test set of Common-ToM to answer yes or no questions. We format our prompt by starting with an instruction (which we keep the same for all prompts), a dialog containing the utterance of interest with five previous and future utterances, and the question. For our GPT-3.5 and GPT-4 experiments, we use the default API hyperparameters. We provide further experimental details, prompts, and hyperparameters, as well as details on unsuccessful chain-of-thought experiments, in Appendix C.
@@ 4.4 Our System: ReCoG
Our Full-Representation approach (ReCoG) creates an explicit representation of the cognitive states of the discourse participants, and then uses rules to answer the questions. We closely follow Markowska et al. (2023) in fine-tuning FLAN-T5 and we similarly use a speaker-based window. Specifically, our model receives as input all utterances preceding and/or following the target event as input until it encounters an utterance from the other speaker.
@@
Belief Prediction We first predict the beliefs of the discourse participants towards a target event at each utterance in the dialog. Our input is as follows:
@@
CG Prediction To predict CG, we use the heuristics-based approach from Markowska et al. (2023). This rule-based approach maps a label for speaker and hearer belief to a CG label.
Yes/No Question Answering Now that we have the belief and CG for both discourse participants, we apply the same heuristics we used to determine the query answers for the gold-annotated case, as described in section 3.
@@ 4.5 Results
Zero-Shot GPT-4 performs the best on all metrics. Mistral-7B-Instruct, despite being only 7B parameters, performs better than the 175B parameter GPT-3.5 in a zero-shot setting. We notice a distinct trend: all models capture first order beliefs the best, and decrease in performance as the order increases.
Fine-tuning Compared to the best zero-shot performance from GPT-4, our results show an increase in total accuracy, second order beliefs, and third order beliefs. However, we perform slightly worse in modeling first order beliefs.
ReCoG ReCoG, which explicitly models beliefs and CG among agents in dialog, outperforms every other system. We see a clear boost in performance in all metrics, and more notably similar results among all orders of belief.
@@ 5 Conclusion
We present a new corpus for testing theory of mind (ToM) capabilities, Common-ToM. Unlike previous ToM corpora, Common-ToM uses naturally occurring linguistic data and is not based on agents perceiving certain information or not. We explicitly model belief and CG to capture ToM in a manner directly inspired by cognitive science literature. While LMs are rapidly evolving and scaling in size and capabilities, they still lack a conceptual understanding of beliefs and CG in dialog. Therefore, it is necessary to explicitly model belief, CG, and cognitive states.
@@
While we present preliminary results on Common-ToM, the first ToM corpus to use natural conversation, the original CG corpus is relatively small, containing only four dialogs. Furthermore, as these dialogs only contain conversations in English, our benchmark does not involve tracking ToM in other languages. While the benchmark tests ToM reasoning in a variety of ways, it is by no means comprehensive. In cognitive science, ToM is divided into cognitive (e.g. beliefs, thoughts) and affective (e.g. emotions, desires) with some evidence for fairly independent processes which operate over these two domains Kalbe et al. (2010). We evaluate the cognitive aspect of belief and plan to tackle additional areas, particularly affective, in future work.
@@ Acknowledgements
This material is based upon work supported in part by the National Science Foundation (NSF) under No. 2125295 (NRT-HDR: Detecting and Addressing Bias in Data, Humans, and Institutions); by funding from the Defense Advanced Research Projects Agency (DARPA) under the CCU (No. HR001120C0037, PR No. HR0011154158, No. HR001122C0034) and INCAS (HR01121C0186, No. HR001120C0037, and PR No. HR0011154158) programs; as well as by the The Intelligence Advanced Research Projects Activity (IARPA) under the HIATUS program (contract 2022-22072200005). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF, DARPA, or IARPA.
@@ Appendix A Corpus Details
We show our train/test split in Table 3. Our corpus builds on Markowska et al. (2023), which uses the English-only LDC CALLHOME corpus as a base.
Dataset Label Count
Train No 2899
Yes 2371
Test No 1139
Yes 965
Table 3: Train/test split of yes and no questions for CommonToM.
For convenience, we summarize the label meanings in Table 4 below.
@@
Table 4: Summary of annotation label meanings as described by Markowska et al. (2023).
@@ Appendix C Experiment Details
We perform a standard fine-tuning approach of FLAN-T5-base model which has 250M parameters. We fine-tune for 12 epochs and use a learning rate of 3e-4. Our system is implemented using the PyTorch [^3] https://pytorch.org/ framework and Python programming language. FLAN-T5 model is captured from the HuggingFace transformers (Wolf et al., 2020). We also used Pandas data analysis library pandas development team (2020).

### v2

@@ Abstract
Evaluating the theory of mind (ToM) capabilities of language models (LMs) has recently received a great deal of attention. However, many existing benchmarks rely on synthetic data, which risks misaligning the resulting experiments with human behavior. We introduce the first ToM dataset based on naturally occurring spoken dialogs, Common-ToM, and show that LMs struggle to demonstrate ToM. We then show that integrating a simple, explicit representation of beliefs improves LM performance on Common-ToM.
[…]
[ backgroundcolor=gray!10, linecolor=gray!60, innertopmargin=0, innerbottommargin=0 ]lstlisting
@@ 1 Introduction
In this paper, we introduce Common-ToM– a question answering benchmark based on naturally-occurring, spoken dialogs in English. [^1] https://github.com/cogstates/common-tom Common-ToM uses the notion of common ground (CG) for evaluating ToM. The CG (Wilkes-Gibbs and Clark, 1992; Stalnaker, 2002) is a set of beliefs mutually shared by all participants in a conversation. Other benchmarks have developed ToM questions based on the Sally-Anne test (Wimmer and Perner, 1983; Baron-Cohen et al., 1985), which checks an understanding of information accessibility to determine beliefs. We observe that when the CG is mismatched between the discourse participants, such as at the time of a question or during the repair of a miscommunication, similarly complex problems for ToM arise.
Our main contributions are: (1) arguing that using synthesized data in evaluating the ToM ability of LMs is not conclusive; (2) releasing a corpus for benchmarking ToM based on naturally occurring spoken conversations; (3) showing that LLMs struggle with our benchmark and a simple explicit architecture performs better.
The paper is organized as follows. First, we review some of the relevant literature (§ 2). We then describe the framework and methods used in creating Common-ToM (§ 3). This is followed by experiments and results in Section 4: human performance, using zero-shot and fine-tuned LMs, and using our approach for building a system that explicitly represents beliefs (§ 4.4). Finally, we discuss our key takeaways and conclusions (§ 5).
@@ 2 Related Work
While quite a few corpora annotate author belief (de Marneffe et al., 2019; Saurí and Pustejovsky, 2009), there are relatively few corpora annotated explicitly for CG. Horton and Gerrig (2016); Soubki et al. (2022) focus on limited aspects of CG. In this paper, we use the comprehensive CG corpus we presented in (Markowska et al., 2023), which we discuss in Section 3.
@@
Though useful, these benchmarks are prone to surface-level cues and spurious correlations which have been exploited by LMs to display illusory ToM (Kosinski, 2023). Some work has been done to produce tools based on human-generated text. Bara et al. (2021), like us, exploit the relationship between CG and ToM. They design an experimental setup that records written dialogs of players exchanging information in MineCraft. More recently, Ma et al. (2023) combine the aforementioned datasets (and more) to create a composite benchmark. We build on this work conceptually by producing a new dataset which, in contrast with the work of Bara et al. (2021), is based on dialogs which are collected independently of our interest in ToM; the dialogs are spontaneous and not guided by any experimental setting; and they are spoken dialogs. We note that co-presence (including telepresence) has been identified as an important (though not required) part of human ToM (Galati and Brennan, 2021), and our corpus allows the study of ToM under this common condition using naturalistic data.
@@ 3 Common-ToM
Framework
In philosophy, CG is sometimes treated as the mutual beliefs between two agents Stalnaker (2002), separate from their cognitive states, while in cognitive science it is common to model CG as the belief of an agent about what they and another agent mutually believe (Brown-Schmidt and Duff, 2016). We adopt the latter definition as it is allows CG to represent scenarios like false belief, which would be impossible under the former. This also makes the relationship between CG and ToM explicit: If I believe a proposition is in the CG with you, it means by definition that I believe that you also believe it is in CG – thus, it entails a ToM assumption.
We also follow cognitive literature in assuming that humans do not have a full representation of their interlocutor’s cognitive state at all times, but instead can make necessary inferences when needed (Horton and Brennan, 2016). Specifically, CG does not mean that all consequences of CG are present at all times, and performing the inferences from CG (high-degree knowledge questions, for example) takes time and may be errorful. Our experiments with human annotators (§ 4) confirm this assumption, but also show that humans are much better than zero-shot LMs because they are in fact modeling CG, and can answer the higher-order belief questions as needed from their models of CG. Farkas and Bruce (2010); Eckardt (2016) propose deterministic heuristic algorithms for inferring CG from text. All this prior work motivates our neuro-symbolic approach (§ 4.4).
Base Corpus
In our previous annotation work (Markowska et al., 2023), we use a corpus of 1,710 turns of dyadic dialog across four conversations from CALLHOME, and annotate them for CG. We extract events (i.e., propositions) evoked by each turn and label the beliefs of each discourse participant (DP) towards those events as certainly true (CT+), possibly true (PS), certainly not true (CT-), or we label the DP as not having a belief (NB). Then we determine for each DP if the speaker has just added the event to their view of the CG (JA), already has the event in the CG (IN), or rejects it from the CG (RT). As we do not annotate CG for events before they are introduced in (Markowska et al., 2023), we use NA to indicate this in this study. Note that the annotation assumes that the CG is not actually shared, but rather independently hypothesized by each DP.  [^2] We refer the reader to Appendix A and (Markowska et al., 2023) for additional details regarding the corpus.
@@
To motivate the query creation process, consider the examples in Table 1. The proposition under question is “B is now not smoking”. The annotations from the CG corpus for this proposition throughout discourse time (which we simply measure in turns) are shown above, and resulting first order (A/B believes) and second order (A/B believes that B/A believes) questions are shown below.
@@
At the time indicated, is it the case that (((A/B believes that)1 B/A believes that)2 A/B believes that)3 it is {certainty} true that {proposition}?
@@
Finally, (3) we determine the answers to the generated queries using a set of rules. For first order questions this is straightforward and we simply check that the belief annotation matches the question, with one caveat. If A believes proposition $p$ is certainly true and the query asks if A believes $p$ is possibly true, we consider the answer to be yes. For higher order queries we must use the CG annotations. To illustrate, consider the 2nd order question whether A believes that B believes some proposition $p$. (1) If the certainty is positive polarity (i.e., certainly or possibly) and A believes the proposition to be CG (i.e., JA or IN), then resolve the answer just as in the first order case described above. (2) Otherwise, if the certainty is negative polarity (i.e. certainly not) and A is aware it was rejected from CG (i.e., RT) and the certainty of the query matches the certainty of B, then label the question yes. (3) For all other cases, label the answer no. The whole set of heuristics is in Appendix E.
Split Answer Count
Train No 2899
Yes 2371
Test No 1139
Yes 965
Table 2: Number of questions in Common-ToM with yes and no answers broken down by split.
Corpus Statistics
This results in 7,374 queries to probe for yes/no answers regarding the beliefs of speakers from CALLHOME. There are a roughly equal number of first, second, and third order queries. The corpus is partitioned using the same splits as was done in (Markowska et al., 2023) – three conversations for training, and a held out fourth conversation for testing. The counts by query answer are shown in Table 2. Additional information regarding the data and splits is in Appendix A.
@@ 4.1 Setup
Data
For fine-tuning experiments we use the train and test splits for Common-ToM as described in Section 3. When prompting, we evaluate only on the test conversation to maintain comparability. For all experiments we do not do any hyperparameter search or hyperparameter tuning.
Random Performance
We perform a random baseline by randomly selecting “Yes” or “No” answers with their probabilities proportional to their respective frequencies in the training set.
Human Performance
We randomly sample sixty queries from our test set with twenty questions from each order. Annotators are two graduate students instructed to answer the queries as best they can. We report accuracy for these sixty questions.
@@ 4.2 Zero-Shot
We format our prompt by starting with an instruction (which we keep the same for all prompts), a dialog containing the utterance of interest with five previous and future utterances, and the question. For our GPT-3.5 and GPT-4 experiments, we use the default API hyperparameters. We provide further experimental details, prompts, and hyperparameters, as well as details on unsuccessful chain-of-thought experiments, in Appendix C.
@@ 4.4 Our System: ReCoG
Our Full-Representation approach (ReCoG) creates an explicit representation of the cognitive states of the discourse participants, and then uses rules to answer the questions. We closely follow our previous work in (Markowska et al., 2023) in fine-tuning FLAN-T5 and we similarly use a speaker-based window. Specifically, our model receives as input all utterances preceding and/or following the target event as input until it encounters an utterance from the other speaker.
@@
Belief Prediction
We first predict the beliefs of the discourse participants towards a target event at each utterance in the dialog. Our input is as follows:
@@
CG Prediction
To predict CG, we use the heuristics-based approach we used in our previous work (Markowska et al., 2023). This rule-based approach maps a label for speaker and hearer belief to a CG label. [^3] See Appendix B for details on this algorithm.
Yes/No Question Answering
Now that we have the belief and CG for both discourse participants, we apply the same heuristics we used to determine the query answers for the gold-annotated case, as described in Section 3.
@@ 4.5 Results
Zero-Shot
GPT-4 performs the best on all metrics. Mistral-7B-Instruct, despite being only 7B parameters, performs better than the 175B parameter GPT-3.5 in a zero-shot setting. We notice a distinct trend: all models capture first order beliefs the best, and decrease in performance as the order increases.
Fine-tuning
Compared to the best zero-shot performance from GPT-4, our results show an increase in total accuracy, second order beliefs, and third order beliefs. However, we perform slightly worse in modeling first order beliefs.
ReCoG
ReCoG, which explicitly models beliefs and CG among agents in dialog, outperforms every other system. We see a clear boost in performance in all metrics, and more notably similar results among all orders of belief.
Discussion
There are some interesting trends in these results. All the LM-only models see a decrease in performance as the order of the query increases, just as humans do, though this decrease is not as large as is observed for humans. This might suggest that the models are in some ways “human-like” in their mistakes. However, upon closer inspection we see further differences.
Since the same proposition is asked about for multiple orders of belief, we can look at how often the model gets all answers for a proposition correct. Even though fine-tuning Mistral only provides a roughly 3 point boost to overall accuracy, this consistency in answers changes dramatically with Mistral, going from getting all questions for a proposition correct 20% of the time in the zero-shot setting to 61% of the time when fine-tuned. Our human sample is too small to perform a similar comparison but Kim et al. (2023) observed human consistency to be in the high 80s for a similar task on their synthetic benchmark.
We can also look at how correctness for first order belief correlates with correctness of higher order beliefs. Again we see fine-tuning makes a larger difference than accuracy suggests. The correlation between per-proposition 1st and 2nd order accuracy goes from $r=0.43$ for the zero-shot model to $r=0.90$ for the fine-tuned model. An even larger difference is observed between 1st and 3rd order ($r=0.20$ to $r=0.93$).
This all suggests that the zero-shot LM behavior is markedly different from the fine-tuned model, and from humans. Even when making correct predictions, they do so using a very different pattern.
@@ 5 Conclusion
We present a new corpus for testing theory of mind (ToM) capabilities, Common-ToM. Unlike previous ToM corpora, Common-ToM uses naturally occurring linguistic data and is not based on agents perceiving certain information or not. We explicitly model belief and CG to capture ToM in a manner directly inspired by cognitive science literature. While LMs are rapidly evolving and scaling in size and capabilities, they still lack a conceptual understanding of beliefs and CG in dialog. Therefore, for an LM to interact naturally it is necessary to explicitly model belief, CG, and cognitive states.
@@
While we present preliminary results on Common-ToM, the first ToM corpus to use natural spoken conversation, the original CG corpus is relatively small, containing only four dialogs. Furthermore, as these dialogs only contain conversations in English, our benchmark does not involve tracking ToM in other languages. While the benchmark tests ToM reasoning in a variety of ways, it is by no means comprehensive. In cognitive science, ToM is divided into cognitive (e.g. beliefs, thoughts) and affective (e.g. emotions, desires) with some evidence for fairly independent processes which operate over these two domains Kalbe et al. (2010). We evaluate the cognitive aspect of belief and plan to tackle additional areas, particularly affective, in future work.
@@ Acknowledgements
This material is based upon work supported in part by the National Science Foundation (NSF) under No. 2125295 (NRT-HDR: Detecting and Addressing Bias in Data, Humans, and Institutions); by funding from the Defense Advanced Research Projects Agency (DARPA) under the CCU (No. HR001120C0037, PR No. HR0011154158, No. HR001122C0034) and INCAS (HR01121C0186, No. HR001120C0037, and PR No. HR0011154158) programs; as well as by the Intelligence Advanced Research Projects Activity (IARPA) under the HIATUS program (contract 2022-22072200005). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF, DARPA, or IARPA.
@@
We thank our ARR reviewers, whose comments have contributed to improving the paper.
@@ Appendix A Corpus Details
Common-ToM builds on (Markowska et al., 2023), which uses the English-only LDC CALLHOME corpus as a base. We show counts for our train/test split in Table 2. The training split contains dialogs 4245, 4248, and 4310 while the test split contains dialog 4431.
For convenience, we summarize the label meanings in Table 4 and the distribution of labels for Common-ToM in Table 5.
@@
NA An event $e$ that has no annotation for CG
because it has not yet been introduced.
Table 4: Summary of annotation label meanings as described in (Markowska et al., 2023).
[…]
CT- CT- RT RT 1,746
CT+ CT+ JA JA 1,632
CT+ CT+ IN IN 1,494
PS PS JA JA 1,152
PS CT+ JA JA 180
PS CT+ NA NA 144
CT- NB NA NA 90
PS PS NA NA 90
NB CT- NA NA 72
PS CT- NA NA 72
CT+ NB NA NA 72
NB CT+ NA NA 54
PS NB NA NA 54
CT+ CT- RT RT 54
CT+ PS JA JA 54
CT- PS NA NA 54
CT- PS RT RT 36
CT+ PS NA NA 36
NB PS NA NA 36
PS CT- RT RT 36
PS PS IN IN 36
CT+ CT+ NA NA 36
CT+ CT+ JA IN 18
NB CT+ JA JA 18
CT- PS NA JA 18
CT+ NB RT RT 18
CT+ CT- IN RT 18
CT- CT+ RT RT 18
CT- CT+ NA NA 18
PS NB RT RT 18
Total 7,374
Table 5: Counts of belief labels included in the corpus.
@@ Appendix C Experiment Details
We perform a standard fine-tuning approach of FLAN-T5-base model which has 250M parameters. We fine-tune for 12 epochs and use a learning rate of 3e-4. Our system is implemented using the PyTorch framework and Python programming language. FLAN-T5 model is captured from the HuggingFace transformers (Wolf et al., 2020). We also used Pandas data analysis library pandas development team (2020).